OpenStudy (anonymous):

Help me Solve the following Definite Integral in detail: \[\int\limits_{1}^{4}\frac{ 3x^3-2x^2+4 }{ x^2}\]

OpenStudy (anonymous):

\[ \int\limits_{1}^{4}\frac{ 3x^3-2x^2+4 }{ x^2}dx=\\ \int\limits_{1}^{4} \left( 3 x -2 - \frac 4 {x^2} \right)dx \] Can you finish it now?

OpenStudy (anonymous):

I can not :(

OpenStudy (anonymous):

\[ \left[ \frac {3 x^2}2 - 2 x - \frac 4 x \right]_1^4 \]

OpenStudy (anonymous):

My first post has a sign misprint. Let me correct ti now\[ \int\limits_{1}^{4}\frac{ 3x^3-2x^2+4 }{ x^2}dx=\\ \int\limits_{1}^{4} \left( 3 x -2 + \frac 4 {x^2} \right)dx\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the result is this last part that you posted?

OpenStudy (anonymous):

No, you have to compute the value of the expression at 4 than at 1 and you subtract the two values

OpenStudy (anonymous):

If we call \[ F(x)=\frac {3 x^2}2 - 2 x - \frac 4 x\\ \] The answer is \[ F(4)- F(1) \]

OpenStudy (anonymous):

\[ F(4)=15\\ F(1)=-\frac 92\\ F(4)-F(1)=\frac {39}2 \]

OpenStudy (anonymous):

= (3/2 . 4² - 2.4 - 4/4) - (3/2.1² - 2.1 - 4/1) =(24 - 8 - 1) - ( 3/2 - 6) = 15 - (-9/2) = 15 + 9/2 = 39/2

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you friend ..

OpenStudy (anonymous):

YW